OpenStudy (anonymous):

Find the unknown length. A=5 b=12

OpenStudy (anonymous):

What?

OpenStudy (danjs):

missing something

OpenStudy (anonymous):

That's all it gave me

OpenStudy (danjs):

the value of two variables. What is the unknown length?, still unknown

OpenStudy (anonymous):

A^2+b^2=c^2

OpenStudy (anonymous):

replace the given A and B to the equation that you have

OpenStudy (anonymous):

A=5 b=12 A^2+b^2=c^2 ( at this time, don't care about the c, we'll deal with it later on) 5^2 +12^2=c^2 (5*5) + (12*12) =c^2 25+144=c^2 169=c^2 Okay, now this time we care about the letter c. When you have C is unknow #, don't worry about it, it normally will be the answer from the sum of A and B. But in this case, it's still a unknow # because of the square! So when It square, you do square roof! like this \[\sqrt{169}=c\] And why don't you square roof the C? because square is square roof so when you square roof the #, aka you pass the square to the number, under square roof form, and why you do that? Search google, I dunno how to explain, so that this step, you can take the # divide for 1 over 2 or use calculator \m/ so c = 13 tada~~~~